# Redshift Drift in Relativistic N-Body Simulations

### v1

@@ 5 Conclusions
The cosmological redshift drift promises to be the first observable that directly constrains the cosmic expansion rate. To prepare these measurements and study the effect the structures in the universe will have on them, we have studied the redshift drift using general relativistic N-Body simulations in a $\Lambda$CDM universe. We have calculated the redshift drift directly from the past light cone, without approximations, for 10 random observers.
Except at very low redshifts $z\lesssim 0.05$, where we see a large scatter due to the local environment, the mean redshift drift follows the expected FLRW background signal with a relative deviation of approximately $0.05\%$ for all observers. For the largest redshifts in our simulation $0.28\lesssim z\lesssim 0.38$, all observers see a smaller redshift drift than the background result. Since the difference is small, this is most likely not relevant for observations.
@@
Lastly, we have calculated angular power spectra for the redshift drift fluctuations and compared them to predictions with linear perturbation theory. On linear scales, we find excellent agreement between the theoretical predictions and our simulation result in both shape and amplitude. This is contrary to the preprint [17] where significant differences between simulation and linear power spectra were found. Additionally, we found a surprisingly large non-linear signal arising from those sources with high peculiar acceleration and velocity, which has not been seen before, although several earlier studies have considered redshift drift fluctuation power spectra.
@@ Acknowledgements.
AO would like to thank Queen Mary University of London for hospitality during part of this project and Asta Heinesen for valuable discussions. We thank Hayley J. Macpherson and Asta Heinesen for permission to reuse data from [13]. We also thank the authors of [28] for sharing their modified version of CLASS and Pedro Bessa for discussing the results. SMK and AO are funded by VILLUM FONDEN, grant VIL53032. The simulations used in this paper were run using the UCloud interactive HPC system managed by the eScience Center at the University of Southern Denmark.
@@ Appendix A Comparison to Bessa et al 2023
In the main text we mention that the angular power spectrum derived in Sect. 3.2 has earlier been derived in [28] in a different manner. The authors of [28] derived the spectrum in a fully gauge-independent manner and included also metric fluctuations, which we neglected here. According to our results in Sect. 4.1 and also the conclusions in [28], neglecting the metric perturbations is very well justified and should not influence the results. Adopting the convention used in [28], the redshift drift fluctuations are represented as
@@
Using the convention by [28], we calculate the redshift drift fluctuation power spectrum once with our result (41) and once with the CLASS modification implemented by [28]. We calculate the spectrum for redshifts $z=0.1,0.5,1,5$ with a Gaussian window function of width $\Delta z=0.01$, the same as in figure 3 of [28]. The result can be seen in Fig. 11. One can see that there is a clear difference in amplitude between the two results, especially for low redshifts. Except the neglected potential perturbations, which [28] showed do not change their power spectrum, our analytical expression for the power spectrum are equivalent. The difference seen in Fig. 11 is therefore quite surprising. Since our results agree well with our simulation results, we are inclined to believe our implementation.
Figure 11: Comparison between predictions for the angular power spectrum of redshift drift fluctuations from perturbation theory as calculated in this paper and the results by [28].

### v2

@@ 6 Conclusions
[…]
Except at very low redshifts, $z\lesssim 0.05$, where we see a large scatter due to the local environment, the mean redshift drift follows the expected FLRW background signal with a relative deviation of approximately $0.05\%$ for all observers. For the largest redshifts in our simulation $0.28\lesssim z\lesssim 0.38$, all observers see a smaller redshift drift than the background result. Since the difference is small, this is most likely not relevant for observations.
@@
Lastly, we have calculated angular power spectra for the redshift drift fluctuations and compared them to predictions from linear perturbation theory. On linear scales, we find excellent agreement between the linear predictions and our simulation result in both shape and amplitude. This is contrary to the preprint [Bessa2024] where significant differences between simulation and linear power spectra were found. Additionally, we found a surprisingly large non-linear signal arising from those sources with high peculiar acceleration and velocity, which has not been seen before, although several earlier studies have considered redshift drift fluctuation power spectra.
@@ Acknowledgements.
AO would like to thank Queen Mary University of London for hospitality during part of this project and Asta Heinesen for valuable discussions. We thank Hayley J. Macpherson and Asta Heinesen for permission to reuse data from [Koksbang2024]. We also thank the authors of [Bessa2023] for sharing their modified version of CLASS and Pedro Bessa for discussing the results. SMK and AO are funded by VILLUM FONDEN, grant VIL53032. The simulations used in this paper were run using the UCloud interactive HPC system managed by the eScience Center at the University of Southern Denmark.
@@ Appendix A Comparison to Bessa et al 2023
In the main text we mention that the angular power spectrum derived in Sect. 3.2 has earlier been derived in [Bessa2023] in a different manner. The authors of [Bessa2023] derived the spectrum in a fully gauge-independent manner and included also metric fluctuations, which we neglected here. According to our results in Sect. 4.1 and also the conclusions in [Bessa2023], neglecting the metric perturbations is very well justified and should not influence the results. Adopting the convention used in [Bessa2023], the redshift drift fluctuations are represented as
@@
Using the convention by [Bessa2023], we calculate the redshift drift fluctuation power spectrum once with our result (41) and once with the CLASS modification implemented by [Bessa2023]. We calculate the spectrum for redshifts $z=0.1,0.5,1,5$ with a Gaussian window function of width $\Delta z=0.01$, the same as in figure 3 of [Bessa2023]. The result can be seen in Fig. 11. One can see that there is a clear difference in amplitude between the two results, especially for low redshifts. Except the neglected potential perturbations, which [Bessa2023] showed do not change their power spectrum, our analytical expression for the power spectrum are equivalent. The difference seen in Fig. 11 is therefore quite surprising. Since our results agree well with our simulation results, we are inclined to believe our implementation.
Figure 11: Comparison between predictions for the angular power spectrum of redshift drift fluctuations from perturbation theory as calculated in this paper and the results by [Bessa2023].